xy dx + (2x ^ 2 + 3y ^ 2 - 20) * d * y = 0 .
1
Expert's answer
2021-09-23T17:30:21-0400
"M(x,y)=xy"

"M_y=x"

"N(x,y)=2x^2+3y^2-20"

"N_x=4x"

"M_y=x\u22604x=N_x"

The equation is not exact.


"xy^4dx+y^3(2x^2+3y^2-20)dy=0"

"M(x,y)=xy^4"

"M_y=4xy^3"

"N(x,y)=y^3(2x^2+3y^2-20)"

"N_x=4xy^3"

"M_y=4xy^3=N_x"

The equation is exact.

There exists a function "f" for which


"\u2202f\/\u2202x=M(x,y)"

We can find "f"  by integrating "M(x,y)" with respect to "x" while holding "y" constant:


"f(x, y)=\int M(x,y)dx+g(y)"

"f(x, y)=\int xy^4dx+g(y)=\dfrac{1}{2}x^2y^4+g(y)"

Taking the partial derivative of the last expression with respect to "y" and setting the result equal to "N(x,y)" gives


"2x^2y^3+g'(y)=2x^2y^3+3y^5-20y^3"

"g'(y)=3y^5-20y^3"

Integrate with respect to "y"


"g(y)=\int(3y^5-20y^3)dy=\dfrac{1}{2}y^6-5y^4-C"

Hence


"f(x, y)=\dfrac{1}{2}x^2y^4+\dfrac{1}{2}y^6-5y^4-C"

The solution of the equation in implicit form is


"\dfrac{1}{2}x^2y^4+\dfrac{1}{2}y^6-5y^4=C"